OpenStudy (anonymous):

Which of the following is equivalent to x5y2/xy2when x does not equal 0 and y does not equal 0 @ganeshie8

OpenStudy (bibby):

division of exponents with the same base = subtraction of the powers \[\frac {x^6} {x^2} = x^4\]

OpenStudy (anonymous):